OpenStudy (anonymous):

The graph of a system of equations with the same slope will have no solutions. Always Sometimes Never

OpenStudy (anonymous):

it will have a solution

OpenStudy (texaschic101):

same slope, different y intercepts = no solutions same slope, same y intercept = infinite solutions so what do you think ?

OpenStudy (anonymous):

sometimes

OpenStudy (texaschic101):

I agree....sometimes

OpenStudy (anonymous):

thnx guys

OpenStudy (texaschic101):

no problem :) glad to help

OpenStudy (anonymous):

Given the equation 4x + 2y = 9, which equation below would cause an inconsistent-independent system? 6x + 12y = 13 2x + y = -6 x - 2y = 7 8x - 4y = -5 can u explain how to do this too?? its confusing

OpenStudy (texaschic101):

well....an inconsistent system would have no solutions...they do not cross. independent system has 1 solution....so I do not understand this...

OpenStudy (anonymous):

alright

OpenStudy (anonymous):

@phi

OpenStudy (anonymous):

what is a consistent tho?

OpenStudy (texaschic101):

Because in an independent system, slopes are different and in an inconsistent system. the slopes are the same.....so I am confused as well

OpenStudy (anonymous):

dp consitents have infinity solutions?

OpenStudy (texaschic101):

consistent has at least 1 solution inconsistent has 0 solutions (parallel line) dependent has infinite solutions (same line) independent has 1 solution

OpenStudy (phi):

** cause an inconsistent-independent system?*** they are asking for two equations that *never intersect* i.e. no solution that happens when they have the same slope but different y intercepts

OpenStudy (texaschic101):

but doesn't independent mean 1 solution.....the question is messing me up when it asks for inconsistent-independent

OpenStudy (anonymous):

ok thank u @phi

OpenStudy (texaschic101):

sorry Milena.....I was confused :(

OpenStudy (anonymous):

its alright

OpenStudy (texaschic101):

the lines are gonna be parallel...do you have the answer yet ?? or do you need help finding it ?

OpenStudy (anonymous):

can u help me?

OpenStudy (texaschic101):

put all the equations in y = mx + b form.....now compare the slopes....you need same slopes but different y intercepts

OpenStudy (anonymous):

ok i will od tht

OpenStudy (anonymous):

it should be the second choice i think

OpenStudy (texaschic101):

4x + 2y = 9 2y = -4x + 9 y = -2x + 9/2 --- slope is -2 and y int is 9/2 6x + 12y = 13 12y = -6x + 13 y = - 6/12 + 13/12 -- slope is -1/2 and y int = 13/12 (nope, not this one) 2x + y = -6 y = -2x - 6 -- slope is -2 and y intercept is -6 (same slope, different y intercepts...THIS ONE

OpenStudy (texaschic101):

yes...it is second choice

OpenStudy (anonymous):

thank u sooo much!!

OpenStudy (texaschic101):

I am glad phi came...because I thought independent meant 1 solution

OpenStudy (anonymous):

oh yeah tht was good